Q3. Ahmed has been offered a 10-year bond issued by Homer, Inc., at a price of $800. The bond has a coupon rate of 7 percent and pays the coupon semiannually. Similar bonds in the market will yield 10 percent today.

1- What should be the price of this bond?

2- Should Ahmed buy the bonds at the offered price?

Please answer it with computer typing.
Answer #1

Hi

Lets assum face value = $1000

Here semiannual coupon of bond C=1000*7%/2 = $35

Semi annual Yield r= 10/2 = 5%

Period t = 10*2 = 20 semiannual

Price of bond = (C/r)*(1-(1+r)^-t) + F/(1+r)^t

   = (35/0.05)*(1-1.05^-20) + 1000/1.05^20

   = 700*0.623111 + 376.89

= $813.07

Since actual price is $800 which is less than calculated price hence Ahmed should buy the bonds.

Thanks